Applications of Mechanics of Differential Equation in Physics


1
Expert's answer
2022-03-28T13:44:25-0400

Differential equation is important in solving mechanical problems

  1. Equation of motion; s=vt,
  2. Schrodinger's Equation(state function equation);"ih\frac{d}{dt}\Psi(r^{\rightarrow},t)=[\frac{-h^2}{2m}]\Delta^2 +V(r^{\rightarrow},t)\Psi(r^{\rightarrow},t)"
  3. Newton's law of cooling;"\frac{dT}{dt}=\alpha(T-T_m)"